OpenStudy (anonymous):

FREE FAN AND MEDAL!!!!!!! Numbers p and q are negative integers. Which statements are always true? Choose all answers that are correct. A. p + q is a negative integer B. p • q is a positive integer C. p – q is a negative integer D. is a negative integer

OpenStudy (anonymous):

@chosenmatt

OpenStudy (chosenmatt):

it would be the first one because

OpenStudy (chosenmatt):

wait

OpenStudy (texaschic101):

A. true -- a neg + a neg = a neg B. true -- a neg times a neg = a positive C. false -- example :(-9 - (-10) = -9 + 10 = 1) D. Can't understand the last one

OpenStudy (chosenmatt):

yeah

OpenStudy (anonymous):

THX BRA

OpenStudy (anonymous):

jk

OpenStudy (chosenmatt):

no problem remember tag me for more help :)

OpenStudy (chosenmatt):

good job @texaschic101

OpenStudy (texaschic101):

thank you...I try